# Fractal ground state of ion chains in periodic potentials

### v1

@@ Abstract
Trapped ions in a periodic potential are a paradigm of a frustrated Wigner crystal. The dynamics is captured by a long-range Frenkel-Kontorova model. The classical ground state can be mapped to the one of an antiferromagnetic spin chain with long-range interactions in a magnetic field, whose strength is determined by the mismatch between chain’s and substrate lattice’s periodicity. The mapping is exact when the substrate potential is a piecewise harmonic potential and holds for any two-body interaction decaying as $1/r^{\alpha}$ with the distance $r$. The ground state is a devil’s staircase of regular, periodic structures as a function of the mismatch, whose range of stability depends also on the coefficient $\alpha$. While the staircase is well defined in the thermodynamic limit for $\alpha>1$, for Coulomb interactions, $\alpha=1$, it disappears and the sliding-to-pinned transitions becomes crossovers. However, due to the logarithmic convergence to the thermodynamic limit characteristic of the Coulomb potential, the staircase is found for any finite number of ions. We discuss the experimental parameters as well as the features that allow one to observe and reveal our predictions in experimental platforms. These dynamics are a showcase of the versatility of trapped ion platforms for exploring the interplay between frustration and interactions.
@@ I Introduction
Interfacing ion chains with optical lattices, as illustrated in Fig. 1, implements a simulator of nanofriction Vanossi et al. (2013). In fact, the Hamiltonian can be reduced to an extended Frenkel-Kontorova (FK) model García-Mata, I. et al. (2007); Pruttivarasin et al. (2011); Cetina et al. (2013); Cormick and Morigi (2013). The FK model describes the interaction of an elastic crystal with an underlying periodic substrate in one dimension Braun and Kivshar (2004). Frustration emerges from the mismatch between the periodicity of the elastic crystal and of the substrate. The ground-state phase diagram of the FK model has been extensively studied for nearest-neighbour interactions: When the corresponding ratio is an incommensurate number, at zero temperature the FK model reproduces the essential features of the stick-slip motion characteristic of static friction, with a continuous transition from sliding to pinning at finite lattice depths. As a function of the mismatch, the ground state is non-analytic and has a form of a devil’s staircase, whose steps correspond to the regime of stability of a commensurate structure, i.e.  a periodic structure pinned by the lattice Aubry and Le Daeron (1983). The transition to a sliding phase is characterized by the proliferation of kinks, namely, of local distributions of excess particles (or holes) in the substrate potential Braun and Kivshar (2004). Realizations with trapped ions observed several features of this dynamics: Stick-slip motion has been reported in experiments with few ions Gangloff et al. (2015); Bylinskii et al. (2015, 2016), pinning by an external lattice has been observed Linnet et al. (2012), the onset of the Aubry transition has been measured in an implementation simulating a deformable substrate Kiethe et al. (2017), and the kinks density has been revealed in small chains as a function of the mismatch Gangloff et al. (2020).
@@
Figure 1: Illustration of a trapped-ion realization of the Frenkel-Kontorova model. A chain of cold ions (orange circles) at uniform distance $d_{0}$ minimizing the Coulomb repulsion. The ions are interacting with the standing wave of a laser field (blue shades), forming a periodic potential with periodicity $a$. For $d_{0}\gg a$ the ions’ motion about the equilibrium positions is described by their harmonic vibrations about the equilibrium configuration and the dynamics is captured by a Frenkel-Kontorova model with long-range elastic forces.
@@
These results show the versatility of trapped ion platforms as quantum simulators. Recent progress in cooling large ion chains Lechner et al. (2016); Feng et al. (2020) and loading ions in optical lattices Schmidt et al. (2018); Hoenig et al. (2023) pave the way towards studying kinks dynamics and their mutual interactions, thus shedding light into the interplay between geometric frustration and quantum fluctuations. In this regime, long-range forces, such as the Coulomb repulsion, qualitatively modify the kinks and the nature of their interactions Pokrovsky and Virosztek (1983). A systematic study can be performed in the continuum limit, when the substrate potential is a small perturbation to the chain’s interaction and the kinks are Sine-Gordon solitons for nearest-neighbor interactions Frank et al. (1949). Then, the long-range interactions modify the Sine-Gordon equation introducing an integral term Landa et al. (2020) and the long-range Sine-Gordon model can be mapped to an extended massive (1+1) Thirring Hamiltonian, where the solitons are charged positive energy excitations over a Dirac sea Menu et al. (2023). This theory has a predictive power for ion chains provided the average interparticle distance is orders of magnitude larger than the potential periodicity, allowing one to discard the discrete nature of the charge density distribution. The theory does not capture the opposite limit, where either the number of ions is limited to few dozens Bylinskii et al. (2016); Kiethe et al. (2018); Gangloff et al. (2020) and/or the depth of the substrate potentials localizes the kinks at chain sizes of a few ions as in Refs. Landa et al. (2013); Partner et al. (2013). In some treatments the discrete nature of the charge distribution can be theoretically described as corrections to the continuum limit Willis et al. (1986); Braun et al. (1990); Chelpanova et al. (2024), leading to effective soliton-phonon collisions Willis et al. (1986).
@@
In the present work, we choose a different approach and start from a discrete distribution of interacting particles. Due to the long-range interactions, the ground state emerging from the competition of interactions and substrate potential cannot be found by means of the ingenious dynamical map of Ref. Aubry and Le Daeron (1983). We instead implement the method of Hubbard Hubbard (1978), and map the ground state configuration of the long-range, Coulomb Frenkel-Kontorova model to the one of an antiferromagnetic spin chain in the presence of a transverse field. The mapping is exact for a periodic substrate composed of piecewise harmonic oscillators Beyeler et al. (1980), illustrated in the upper panel of Fig. 2, and is amenable to analytical solutions. Despite the theoretical abstraction, we show that this mapping sheds light on the properties of realistic substrate potentials, such as an optical lattice.
@@
The presentation of our study is organised as follows. In Sec. II it is shown that the ground state and low-energy excitations of a Wigner crystal of ions in a linear Paul trap are described by a Frenkel-Kontorova (FK) model where the oscillators of the elastic crystal interact via the long-range Coulomb interactions. This Section reviews the arguments presented in Refs. García-Mata, I. et al. (2007); Pruttivarasin et al. (2011); Cormick and Morigi (2013) and sets the stage for our analysis. The ground state is determined in Sec. III within a mean-field approach, which discards the kinetic energy. Here, we assume a specific function of the periodic substrate and map the continuous-variable problem onto an antiferromagnetic Ising model with long-range interactions and in the presence of a magnetic field, as illustrated in the lower panel of Fig. 2. Our mapping extends the study of Ref. Beyeler et al. (1980) to a Wigner crystal and allows us to show that the ground state is a devil’s staircase as a function of the mismatch between the lattice periodicity and the characteristic interparticle distance. It allows us, moreover, to determine the interval of stability of the individual commensurate structures as a function of the temperature. In Sec. IV we determine the low-energy excitations of an ion chain in a periodic potential across the Aubry transition and identify the experimental signatures. We then discuss the order of magnitude of quantum fluctuations by means of a semiclassical ansatz. The conclusions are drawn in Sec. V, where we provide an outlook of the directions of studies that our study opens towards the systematic characterization of the interplay between long-range interactions and geometric frustration with cold atoms platforms.
@@ II Chain of interacting particles in a periodic potential
This section reviews the basic assumptions and the steps that connect the Hamiltonian of a one-dimensional Wigner crystal of ions in a periodic potential with a Frenkel-Kontorova model of oscillators interacting via power-law decaying forces. We then generally discuss the geometric properties of the ground state using the characterization of Hubbard Hubbard (1978) and introduce the quantities that will be important for performing the mapping in Sec. III. We refer the interested reader to Refs. García-Mata, I. et al. (2007); Pruttivarasin et al. (2011), where it was proposed to study the sliding-to-pinning transition using Wigner crystals of trapped ions in optical lattices.
@@ II.1 Extended Frenkel-Kontorova model
In order to link the model of Eq. (1) with the paradigmatic Frenkel-Kontorova model, we assume that the particles are localized about the equilibrium positions of the potential $W(x)$, and perform a Taylor expansion of the interaction $W(x)$ about the classical equilibrium positions $x_{j}^{(0)}$ assuming that the average interparticle distance $d_{0}=L/N$ is much larger than the lattice periodicity $a$, thus $x_{j}^{(0)}=jd_{0}$. We denote by $u_{j}$ the local displacement of the particles $u_{j}$ from $x_{j}^{(0)}$, such that $x_{j}=x_{j}^{(0)}+u_{j}$. In second order in the expansion in the small parameter $u_{j}/d$ the potential reads:
@@ II.2 Potential of the vacant sites
Equation (II.2) is the starting point for performing a mapping to a potential of interacting spins. A crucial part of this mapping consists of eliminating the displacement variables $\delta x_{j}$ and rewriting the potential energy only in terms of the vacant-site variables $h_{n}$, which in turn will be mapped onto spins.
@@ III Ground state of the piece-wise parabolic potential
Figure 3: Coefficients of the interacting potential for the vacant sites. The coefficient $J_{\alpha}(r)$, Eq. (LABEL:eq:J:0) are displayed as a function of the distance $r$ (in units of $a$) for a deep ($g=1/10$) and a shallow ($g=3$) substrate lattice. The colors refer to different values of the power-law exponent $\alpha$, see legenda. The logarithmic scale emphasizes the two-fold behavior of the coefficients, which at short distances decay exponentially while at long distances decay algebraically as $1/r^{\alpha}$ for finite exponents $\alpha$. The nearest-neighbor case ($\alpha\to\infty$) is solely characterized by the exponential decay.
We now show that the model of Eq. (II.2) can be mapped to a long-range antiferromagnetic Ising model in the presence of a magnetic field, as illustrated in Fig. 2. The classical ground state of this model is a devil’s staircase as a function of the density $\rho$ Bak and Bruinsma (1982). The mapping we perform is exact when the periodic substrate potential is a sequence of piecewise-truncated parabolas of the form
@@
for $|\delta x|\leq a/2$, see Fig. 2 (a). This functional behavior was assumed by Frenkel and Kontorova in their original model Frenkel and Kontorova (1938), and has been used in several analyses (see, e.g., Aubry (1983); Beyeler et al. (1980); Krajewski and Müser (2005)).
@@ III.1 Mean-field configuration in Fourier representation
For nearest-neighbor interactions ($\alpha\to\infty$) the parameter $g$ controls the transition from sliding, where solitons proliferate ($g\gg 1$), to pinning ($g\ll 1$), where the formation of the soliton is energetically costly.
@@ III.2 The potential for the vacant sites
It is instructive to analyse the generic behavior of the coefficients $J_{\alpha}(r)$ as a function of $r$ for finite power-law exponents $\alpha$, see Fig. 3. We first note that $|\phi_{\alpha}(q)|\leq 2\zeta(\alpha+2)$, with $\zeta(\alpha)$ the Riemann zeta function. By means of an analytic continuation, it becomes visible that the pole of the function $\mathcal{F}_{\alpha}(q)=1+g\phi_{\alpha}(q)$ determines an exponentially decaying behavior with a characteristic length that is monotonically increasing with $g$, see Appendix A. For nearest-neighbor interactions ($\alpha\to+\infty$), the coefficient decays exponentially with a damping length monotonically increasing with $g$. Instead, for finite values of $\alpha$, we observe a two-fold behaviour of $J_{\alpha}(r)$: at short distances the coefficient decays exponentially, whereas at long distances $|r|\gg a$ the coefficient exhibits a power-law tail solely determined by the long-range interactions. In Appendix A we show that the power-law tail takes the form
@@ III.3 The dislocation
III.4 A long-range antiferromagnetic spin model
@@
where $J(r)=J_{\alpha}(r)$. This Hamiltonian describes an Ising antiferromagnetic chain in the magnetic field of strength
@@
The magnetic field, in turn, is proportional to the discommensuration and tends to align the spin, competing with the antiferromagnetic order imposed by the interactions.
@@ III.5 Devil’s staircase
Figure 5: Devil’s staircase of commensurate structures as a function of the mismatch. The plot shows the magnetization, here represented in terms of the ratio $\overline{h}=m/n$ of up-oriented spins, as a function of the magnetic field, here given by the discommensuration $\delta$. The staircase has been numerically determined using the method of Ref. Koziol et al. (2023) for a chain of $N=200$ ions ($\alpha=1$) and $g\ll 1$. The displayed plateaus correspond here to the ratios $m/n$ with $n\leq 20$.
The phase diagram of the spin model of Eq. (23) entails the so-called Aubry transition as well as the commensurate-incommensurate transition Bak (1982). At both transitions the ground state becomes non-analytic. The fractal nature of the ground state becomes visible when considering the so-called magnetization $M$ as a function of the magnetic field. $M$ is given by
$M=\sum S_{j}/N=2\bar{h}-1\,,$ (25)
with $\bar{h}=\langle h\rangle-n_{0}$ the effective discommensuration. In the absence of the substrate potential, $\bar{h}=\delta$, and the magnetization is proportional to the magnetic field. At finite substrate depths instead, $\bar{h}$ exhibits a devil staircase as a function of $\delta$ as shown in Fig. 5. The phase diagram of some commensurate phases in the $g-\delta$ plane is displayed in Fig. 6(a) for the Wigner crystal. One observes that the size decreases as the lattice depth decreases (corresponding to increasing $g$).
Figure 6: (a) Magnetization, $\overline{h}=m/n$, as a function of the discommensuration $\delta$ and of the ratio $g$ between the stiffness and the potential depth. Colored areas represent plateaus corresponding to ratios $\overline{h}$ with $n\leq 8$. In the grey areas the plateaus of the staircase are below the chosen resolution. (b) The magnetization $\overline{h}=m/n$ as a function of the discommensuration $\delta$ and of the temperature $T$ for $g\ll 1$. Colored areas represent plateaus corresponding to ratios $\overline{h}$ with $n\leq 8$. White areas correspond to a paramagnetic, disordered phase. All calculations are performed for a chain of $N=50$ ions interacting via Coulomb repulsion ($\alpha=1$).

### v2

# Fractal ground state of mesoscopic ion chains in periodic potentials

@@ Abstract
Trapped ions in a periodic potential are a paradigm of a frustrated Wigner crystal. The dynamics is captured by a long-range Frenkel-Kontorova model. We show that the classical ground state can be mapped to the one of a long-range Ising spin chain in a magnetic field, whose strength is determined by the mismatch between chain’s and substrate lattice’s periodicity. The mapping is exact when the substrate potential is a piecewise harmonic potential and holds for any two-body interaction decaying as $1/r^{\alpha}$ with the distance $r$. The ground state is a devil’s staircase of regular, periodic structures as a function of the mismatch and of the interaction exponent $\alpha$. While the staircase is well defined in the thermodynamic limit for $\alpha>1$, for Coulomb interactions, $\alpha=1$, we argue that it disappears and the sliding-to-pinned transition becomes a crossover, with a convergence to the thermodynamic limit scaling logarithmically with the chain’s size. Due to this slow convergence, fractal properties can be observed even in chains of hundreds of ions at laser cooling temperatures. These dynamics are a showcase of the versatility of trapped ion platforms for exploring the interplay between frustration and interactions.
@@ I Introduction
Figure 1: Illustration of a trapped-ion realization of the Frenkel-Kontorova model. In the absence of other external potentials, cold ions (orange circles) confined on a line form a chain at uniform distance $d_{0}$ minimizing the Coulomb repulsion. Geometric frustration is introduced by the standing wave of a laser field (blue shades), forming a periodic potential with periodicity $a$. For $d_{0}\gg a$ the ions’ motion about the equilibrium positions is described by their harmonic vibrations about the equilibrium configuration. The dynamics is captured by a Frenkel-Kontorova model with long-range elastic forces.
Interfacing ion chains with optical lattices, as illustrated in Fig. 1, implements a simulator of nanofriction Vanossi et al. (2013). In fact, the Hamiltonian can be reduced to an extended Frenkel-Kontorova (FK) model García-Mata et al. (2007); Pruttivarasin et al. (2011); Cetina et al. (2013); Cormick and Morigi (2013). The FK model describes the interaction of an elastic crystal with an underlying periodic substrate in one dimension Braun and Kivshar (2004). Frustration emerges from the mismatch between the periodicity of the elastic crystal and of the substrate. The ground-state phase diagram of the FK model has been extensively studied for nearest-neighbour interactions: When the corresponding ratio is an incommensurate number, at zero temperature the FK model reproduces the essential features of the stick-slip motion characteristic of static friction, with a continuous transition from sliding to pinning at finite lattice depths. As a function of the mismatch, the ground state is non-analytic and has a form of a devil’s staircase, whose steps correspond to the regime of stability of a commensurate structure, i.e.  a periodic structure pinned by the lattice Aubry and Le Daeron (1983). The transition to a sliding phase is characterized by the proliferation of kinks, namely, of local distributions of excess particles (or holes) in the substrate potential Braun and Kivshar (2004). Experiments with trapped ions observed several features of this dynamics: Stick-slip motion has been reported in chains of few ions Gangloff et al. (2015); Bylinskii et al. (2015, 2016), pinning by an external lattice has been observed Linnet et al. (2012), the onset of the Aubry transition has been measured in an implementation simulating a deformable substrate Kiethe et al. (2017), and the kinks density has been revealed in small chains as a function of the mismatch Gangloff et al. (2020).
@@
These results show the versatility of trapped ion platforms as quantum simulators. Recent progress in cooling large ion chains Lechner et al. (2016); Feng et al. (2020) and loading ions in optical lattices Schmidt et al. (2018); Hoenig et al. (2023) pave the way towards studying kinks dynamics and their mutual interactions, thus shedding light into the interplay between geometric frustration and quantum fluctuations. In this regime, long-range forces, such as the Coulomb repulsion, qualitatively modify the kinks and the nature of their interactions Pokrovsky and Virosztek (1983). A systematic study can be performed in the continuum limit, when the substrate potential is a small perturbation to the chain’s interaction and the kinks are sine-Gordon solitons for nearest-neighbor interactions Frank et al. (1949). Then, the long-range interactions modify the sine-Gordon equation introducing an integral term Landa et al. (2020) and the long-range sine-Gordon model can be mapped to an extended massive (1+1) Thirring Hamiltonian, where the solitons are charged positive-energy excitations over a Dirac sea Menu et al. (2024). This theory has a predictive power for ion chains provided the kink’s size is orders of magnitude larger than the interparticle distance, allowing one to discard the discrete nature of the charge density distribution. The theory does not capture the opposite limit, where either the number of ions is limited to few dozens Bylinskii et al. (2016); Kiethe et al. (2018); Gangloff et al. (2020) and/or the depth of the substrate potentials localizes the kinks in chains composed by few ions as in Refs. Landa et al. (2013); Partner et al. (2013). In some treatments the discrete nature of the charge distribution can be theoretically described as corrections to the continuum limit Willis et al. (1986); Braun et al. (1990); Chelpanova et al. (2024), leading to effective soliton-phonon collisions Willis et al. (1986).
@@
In the present work, we choose a different approach and start from a discrete distribution of interacting particles. Due to the long-range interactions, the ground state emerging from the competition of interactions and substrate potential cannot be found by means of the ingenious dynamical map of Ref. Aubry and Le Daeron (1983). We instead implement the method of Hubbard Hubbard (1978), and map the ground state configuration of the long-range, Coulomb Frenkel-Kontorova model to the one of an antiferromagnetic spin chain in the presence of a magnetic field. The mapping is exact for a periodic substrate composed of piecewise harmonic oscillators Beyeler et al. (1980), illustrated in the upper panel of Fig. 2, and is amenable to analytical solutions. Despite the theoretical abstraction, we show that this mapping sheds light on the properties of realistic substrate potentials, such as an optical lattice.
@@
The presentation of our study is organised as follows. In Sec. II it is shown that the ground state and low-energy excitations of a Wigner crystal of ions in a linear Paul trap are described by a Frenkel-Kontorova (FK) model where the oscillators of the elastic crystal interact via the long-range Coulomb interactions. This Section reviews the arguments presented in Refs. García-Mata et al. (2007); Pruttivarasin et al. (2011); Cormick and Morigi (2013) and sets the stage for our analysis. The ground state is determined in Sec. III within a mean-field approach, which discards the kinetic energy. Here, we assume a specific function of the periodic substrate and map the continuous-variable problem onto an long-range Ising model and in the presence of a magnetic field, as illustrated in the lower panel of Fig. 2. Our mapping extends the study of Ref. Beyeler et al. (1980) to a Wigner crystal and allows us to show that the ground state is a devil’s staircase as a function of the mismatch between the lattice periodicity and the characteristic interparticle distance. It allows us, moreover, to determine the interval of stability of the individual commensurate structures as a function of the temperature. In Sec. IV we determine the low-energy excitations of an ion chain in a sinusoidal potential across the Aubry transition and identify its experimental signatures. We then discuss the order of magnitude of quantum fluctuations by means of a semiclassical ansatz. The conclusions are drawn in Sec. V, where we provide an outlook of the directions of studies that our work opens towards the systematic characterization of the interplay between long-range interactions and geometric frustration with cold atoms platforms.
@@ II Chain of interacting particles in a periodic potential
This section reviews the basic assumptions and the steps that connect the Hamiltonian of a one-dimensional Wigner crystal of ions in a periodic potential with a Frenkel-Kontorova model of oscillators interacting via power-law decaying forces. We then generally discuss the geometric properties of the ground state using the characterization of Hubbard Hubbard (1978) and introduce the quantities that will be important for performing the mapping in Sec. III. We refer the interested reader to Refs. García-Mata et al. (2007); Pruttivarasin et al. (2011), where it was proposed to study the sliding-to-pinning transition using Wigner crystals of trapped ions in optical lattices.
@@ II.1 Extended Frenkel-Kontorova model
In order to link the model of Eq. (1) with the paradigmatic Frenkel-Kontorova model, we assume that the particles are localized about the equilibrium positions of the potential $W(x)$. We perform a Taylor expansion of the interaction $W(x)$ about the classical equilibrium positions $x_{j}^{(0)}$ assuming that the average interparticle distance $d_{0}=L/N$ is much larger than the lattice periodicity $a$, thus $x_{j}^{(0)}=jd_{0}$. We denote by $u_{j}$ the local displacement of the particle $j$ from the equilibrium position $x_{j}^{(0)}$, such that $x_{j}=x_{j}^{(0)}+u_{j}$. In second order in the expansion in the small parameter $u_{j}$ ($u_{j}\ll d_{0}$) the potential reads:
@@ II.2 Potential of the vacant sites
Equation (II.2) is the starting point for performing a mapping to a potential of interacting spins. A crucial part of this mapping consists in eliminating the displacement variables $\delta x_{j}$ and rewriting the potential energy only in terms of the vacant-site variables $h_{n}$, which in turn will be mapped onto spins.
@@ III Ground state of the piece-wise parabolic potential
Figure 3: Scaling of the coefficients of the interacting potential, Eq. (17), with the distance $r$. The coefficient $J_{\alpha}(r)$, Eq. (LABEL:eq:J:0) are displayed as a function of $r$ (in units of $a$) for a deep ($g=1/10$) and a shallow ($g=3$) substrate lattice. The colors refer to different values of the power-law exponent $\alpha$, see legenda. The logarithmic scale emphasizes the two-fold behavior of the coefficients, which at short distances decay exponentially while at long distances decay algebraically as $1/r^{\alpha}$ for finite exponents $\alpha$. The nearest-neighbor case ($\alpha\to\infty$) is solely characterized by the exponential decay.
We now show that the model of Eq. (II.2) can be mapped onto a long-range Ising model in the presence of a magnetic field, as illustrated in Fig. 2(b). The classical ground state of this model is a devil’s staircase as a function of the density $\rho$ Bak and Bruinsma (1982). The mapping we perform is exact when the periodic substrate potential is a sequence of piecewise-truncated parabolas of the form
@@
for $|\delta x|\leq a/2$, see Fig. 2 (a). This functional behavior has been used in several analyses (see, e.g., Aubry (1983); Beyeler et al. (1980); Krajewski and Müser (2005)).
@@ III.1 Mean-field configuration in Fourier representation
For nearest-neighbor interactions ($\alpha\to\infty$) the parameter $g$ controls the transition from sliding, where kinks proliferate ($g\gg 1$), to pinning ($g\ll 1$), where the formation of solitons is energetically costly.
@@ III.2 The potential for the vacant sites
It is instructive to analyse the generic behavior of the coefficients $J_{\alpha}(r)$ as a function of $r$ for finite power-law exponents $\alpha$, see Fig. 3. We first note that $|\phi_{\alpha}(q)|\leq 2\zeta(\alpha+2)$, with $\zeta(\alpha)$ the Riemann zeta function. By means of an analytic continuation, it becomes visible that the pole of the function $\mathcal{F}_{\alpha}(q)=1+g\phi_{\alpha}(q)$ determines an exponentially decaying behavior with a characteristic length that is monotonically increasing with $g$, see Appendix A. For nearest-neighbor interactions ($\alpha\to+\infty$), the coefficient decays exponentially with a damping length monotonically increasing with $g$. Instead, for finite values of $\alpha$, we observe a two-fold behaviour of $J_{\alpha}(r)$: at short distances the coefficient decays exponentially, with a characteristic length depending on $g$, whereas at long distances $|r|\gg a$ the coefficient exhibits a power-law tail solely determined by the long-range interactions. In Appendix A we show that the power-law tail takes the form
@@ III.3 The dislocation
III.4 A long-range Ising model
@@
where $J(r)=J_{\alpha}(r)$. The first term of this Hamiltonian describes an antiferromagnetic Ising chain in a polarizing magnetic field of strength $\sum_{r>0}J(r)$. This polarizing field is shifted by an offset
@@
This additional magnetic field, in turn, is proportional to the discommensuration and tends to align the spin, competing with the antiferromagnetic order imposed by the interactions.
@@ III.5 Devil’s staircase
Figure 5: Devil’s staircase of commensurate structures as a function of the mismatch. The plot shows the magnetization, here represented in terms of the ratio $\overline{h}=m/n$ of up-oriented spins, as a function of the magnetic field, here given by the discommensuration $\delta$. The staircase has been numerically determined using the method of Ref. Koziol et al. (2023) for a chain of $N=200$ ions ($\alpha=1$) and $g=1/10$. The displayed plateaus correspond here to the ratios $m/n$ with $n\leq 20$.
The parameters $\delta$ and $g$ fully characterize the classical ground state. For nearest-neighbour interactions, the phase diagram of the spin model of Eq. (23) entails the so-called Aubry transition, where, at fixed mismatch $\delta$, a critical depth of the potential $V_{0}$ separates the sliding phase from the pinned phase. Here, the spectrum is gapped and the phase is dynamically characterized by stick-slip events. At constant $g$, the phase diagram also entails the so-called commensurate-incommensurate transition: here a critical value of the mismatch separates an ordered (commensurate) phase from the sliding phase where kinks proliferate Bak (1982). At both transitions the ground state becomes non-analytic. The fractal nature of the ground state becomes visible when considering the so-called magnetization $\overline{m}$ as a function of the magnetic field (and thus of the mismatch). The magnetization $\overline{m}$ is defined as
$\overline{m}=\sum S_{j}/N=2\bar{h}-1\,,$ (25)
with $\bar{h}=\langle h\rangle-n_{0}$ the effective discommensuration. In the absence of the substrate potential, $\bar{h}=\delta$, and the magnetization is proportional to the magnetic field. At finite substrate depths instead, $\bar{h}$ exhibits a devil staircase as a function of $\delta$. The staircase exists for all values of $\alpha>1$ Bak and Bruinsma (1982). The Coulomb case $\alpha=1$ is discussed in the next section.
Figure 6: (a) Magnetization, $\overline{h}=m/n$, as a function of the discommensuration $\delta$ and of the ratio $g$ between the stiffness and the potential depth. Colored areas represent plateaus corresponding to ratios $\overline{h}$ with $n\leq 8$. In the grey shaded areas the plateaus of the staircase are below the chosen resolution. (b) The magnetization $\overline{h}=m/n$ as a function of the discommensuration $\delta$ and of the temperature $T$ for $g=1/10$. Colored areas represent plateaus corresponding to ratios $\overline{h}$ with $n\leq 8$. White areas correspond to a paramagnetic, disordered phase. All calculations are performed for a chain of $N=200$ ions interacting via Coulomb repulsion ($\alpha=1$).